### v1

# Design and Verification of the JUNO Liquid Filling Control System

Abstract
Jiangmen Underground Neutrino Observatory (JUNO) is a large-scale neutrino experiment with multiple physics goals including neutrino mass hierarchy, accurate measurement of neutrino oscillation parameters, neutrino detection from supernova, sun, and earth, etc. The Central Detector (CD) of JUNO using 20-kiloton Liquid Scintillator (LS) as target mass with 3% energy resolution at 1 MeV and low radioactive background. To achieve LS filling smoothly and detector running safely, JUNO’s liquid Filling, Overflowing, and Circulating System (FOC) features a control system based on a Programmable Logic Controller and equipped with high-reliable sensors and actuators. This paper describes the design of the FOC automatic monitoring and control system including hardware and software. The control logic including Proportional-Integral-Derivative control, sequential control, and safety interlocks ensures precise regulation of key parameters such as flow rate, liquid level, and pressure. Test results demonstrate the FOC system can satisfy the requirement of JUNO detector filling and running operations.
1 Introduction
The Jiangmen Underground Neutrino Observatory (JUNO) [1] is the next generation neutrino experiment, located about 53 km from the Taishan and Yangjiang nuclear power plants in south China to optimize the precise measurement of the reactor neutrino oscillation parameters. Its primary goal is to determine the neutrino mass ordering - a fundamental question with implications for neutrino mass origins and physics beyond the Standard Model - while also measuring oscillation parameters ($\theta_{12}$, $\Delta m_{21}^{2}$, $|\Delta m_{32}^{2}|$) with per-mil precision. JUNO can also advance the research associated with solar, atmospheric, supernova neutrinos, geoneutrinos, proton decay, etc.
The JUNO Central Detector (CD) is the largest liquid scintillator (LS) detector with 20-kiloton LS contained by a 35.4 meter diameter acrylic sphere, achieving 3% energy resolution at 1 MeV. The acrylic sphere and photo-multiplier tubes (PMT) are supported by a 41.1 meter in diameter stainless steel shell [2]. The whole detector is submerged in 40-kiloton pure water (PW) in water pool (WP) for radiation shielding and active cosmic-ray muon tagging.
[…]
The filling, overflowing, and circulating (FOC) system is in charge of LS filling, overflowing or refilling, and online circulation for LS re-purification [2]. Filling the CD with purified LS is the final step in the detector construction. The process must preserve the structural integrity of the detector and the chemical and optical properties of the LS. During detector filling and data taking, the system should also be able to maintain LS pressure and control the LS level for detector safety. To meet the above requirements, a FOC system is designed with an operation lifetime of 20 years.
The FOC system consists primarily of one storage tank as a buffer container for LS filling and circulation, and two overflow tanks compensating for LS volume fluctuations caused by temperature changes during detector operation. The system also includes pipes, top and bottom chimneys, nitrogen gas system preventing the LS from exposure to oxygen, moisture, and radon. It also features a control system to implement the above functions of FOC.The corresponding Piping and Instrument Diagram (P&ID) of the FOC system is shown in Figure 1.
Figure 1: Primary P&ID of the FOC system
This paper focuses on the control system of the FOC system. The paper is organized as follows. In Section 2, we describe the design, specifications, and principles related to the FOC control system. In Section 3, we describe the principles of control logic and their application scenarios. We present the test performances of the control system in Section 4.
2 System requirement and design
2.1 System requirement
The filling process consists of: (1) synchronous pure water injection into both the CD and WP, (2) replacing pure water in the CD with LS. During the data-taking period, the system must maintain stable CD pressure while limiting liquid level fluctuations to within 20 cm (overflow and refilling condition), and support online LS circulation. To ensure these functions, the FOC control system must meet several major requirements: precision, robustness, scalability, stability, anti-interference, safety and long time data traceability.
Firstly, the FOC control system must exhibit extremely high accuracy. The measurement deviations of all sensors shall be strictly constrained within the allowed range permitted by the process to ensure that every parameter during the filling and detector running accurately reflects the system’s status. High-precision sensors provide a solid foundation for precise system control, with specific accuracy requirements such as a liquid level sensor measurement accuracy $\leq 0.2\%$ full scale (FS) and a flow sensor accuracy reaching ±$\pm 0.1\%$ FS. The total control shall maintain a deviation $\leq 1\%$ between the setpoint and feedback value in a closed-loop control. Pumps shall maintain rotational speed stability within ±$\pm 0.2\%$ of setpoint with $\leq 10$ second standby activation. Regulating valves require ±$\pm 1\%$ positioning accuracy, while on-off valves must achieve < 1 second response time and $<10^{-6}\,\text{mbar}\cdot\text{L}\cdot\text{s}^{-1}$ leak rate.
Secondly, the FOC control system must possess security and scalability. Specifically, it requires safety interlock mechanisms to mitigate risks to detectors during abnormal conditions. A physical emergency stop function is needed to ensure process termination even during system failures. For safety, we also use a "dual-confirmation" protocol to prevent hazardous operations caused by mistakes. Considering potential changes and upgrades in FOC system, the control system must feature modular expandability to avoid system redesigns.
In the end, the FOC control system requires high stability, robustness and anti-interference capability. It must maintain precise control during prolonged continuous operation to avoid risks caused by control deviations and equipment failures. The system should be able to handle environmental fluctuations (e.g., electromagnetic interference, temperature variations, mechanical vibrations) or internal signal noise.
2.2 System design
Based on the above requirements, the layout of the FOC control system is shown in Figure 2. The system integrates 35 sensors, 20 control valves, and 8 pump units, forming a sophisticated control network. In this configuration, the system manages more than 80 sets of analog input/output (I/O) control quantities and more than 150 digital I/O control and alarm signals. Therefore, it mainly consists of 4 layers: the sensor layer, the controller layer, the actuator layer, as well as the alarm and data storage layer.
Figure 2: Layout of FOC control system
2.2.1 Sensor layer
As the critical monitoring terminal, all sensors selected for the FOC system (including level, pressure, temperature, flow rate, etc.) are sourced from Endress+Hauser (E+H) [7], renowned for their stability and reliability with an accuracy better than 0.2% FS. The instrument selection has undergone multiple rounds of demonstration, aligning with the system pipeline or tanks design and meeting the corresponding filling index requirements. For instance, the flowmeters were specifically chosen to match the LS filling rate of $7\ \text{m}^{3}/\text{h}$ with optimal measuring range and precision. These devices have undergone rigorous acceptance testing, including factory tests, offline tests, and online tests. All sensors come pre-equipped with functions such as preset alarm thresholds, alarm signal output, and fault diagnostics. The surface roughness of sensor interfaces in contact with LS meets design requirements and maintains sufficient cleanliness.
Table 1: Main sensor selection for control system
Sensor type Quantity Accuracy (FS) Purpose
Electronic remote differential pressure level gauge 4 0.20% Measure CD liquid level (vertical tube & horizontal tube)
Laser level gauge 1 0.10% Measure CD liquid level
Differential pressure level gauge 2 0.20% Measure liquid level in the upper chimney
Static pressure liquid level gauge 5 0.20% Measure WP liquid level (bottom & surface)
Differential pressure level gauge 5 0.20% Measure tank liquid level
Radar wave level gauge 3 0.10% Measure tank liquid level
Temperature sensor 3 0.20% Measure LS temperature inside tanks
Coriolis mass flowmeter 2 0.10% Measure flow of exchanging LS/water
Pressure transmitter 3 0.15% Measure tank gas pressure
Pressure transmitter 3 0.15% Measure pump discharge pressure
Pressure transmitter 1 0.20% Measure pressure of calibration house
Double-flange pressure transmitter 1 0.10% Measure pressure difference before and after the filter
Vortex flowmeter 2 0.20% Measure water flow during water filling period
The sensor specifications are summarized in Table 1, which aligns with the system’s requirements. Specifically, redundant sensors are implemented at the critical points: (1) CD level monitoring uses 2 pairs electronic remote differential pressure gauges, 1 laser gauge, and 2 chimney differential pressure gauges; (2) WP employs 5 static pressure gauges at different heights; (3) tanks feature dual level sensors; (4) nitrogen pressure is measured by 3 sensors at 3 connected tanks; (5) four temperature sensors are distributed throughout LS pipelines.
These redundant instruments monitor the same parameter simultaneously, with primary and backup devices clearly defined. When the primary instrument fails, the standby device automatically switches to active mode within seconds, leveraging real-time diagnostics and failover algorithms. This hot standby mechanism ensures continuous and accurate data collection, effectively reducing measurement errors and system risks caused by single-instrument failures.
The sensor layer incorporates comprehensive anti-interference measurement to ensure signal reliability. All analog signal transmissions employ single-end grounded shielded twisted-pair cables to minimize electromagnetic interference, while LC filter modules are connected in series with the sensor power supply to suppress power ripple. Critical analog signals undergo through barrier isolators to prevent ground loops. These hardware protection measures work in tandem with the software filtering strategies described in Section 3.1, collectively achieving the goal of signal interference reduction.
2.2.2 Controller layer
The Siemens S7-300 series Programmable Logic Controller (PLC) is chosen as the core component of the FOC control system. Equipped with a high-performance CPU, it has powerful computing and processing capabilities. It can analyze and judge various input signals in micro-second and then quickly issue control commands including equipment start-stop operations and process parameter adjustments. Its modular design enables flexible configuration to meet diverse control requirements while ensuring system reliability through independent module replacement, maintaining uninterrupted operation.
The modular configuration is chosen for FOC PLC as follows: one CPU module for computation, one interface module (IM) for data exchange, six 8-channel analog input (AI) modules for sensor signals, two 8-channel analog output (AO) modules for control, and three digital modules (two 32-channel DI and one 32-channel DO) for switch/alarm handling.
Control software was developed based on Totally Integrated Automation (TIA) Portal [8], which integrates digital planning, operation management, diagnostic capabilities and comprehensive monitoring of the FOC system. The system can acquire and display various parameters in real-time. In addition, we can operate various devices through the control interface, such as adjusting valve openings and controlling pump frequency.
For real-time interaction among different control systems in JUNO, another set of monitoring and control programs is developed based on the Experimental Physics and Industrial Control System (EPICS) [9] framework, which provide functions such as comprehensive monitoring, historical query, and data caching. Moreover, the heartbeat packet program is used to monitor the operational status of PLC. The human-machine interface (HMI) developed based on TIA Portal is shown in Figure 3.
Figure 3: HMI based on TIA Portal
2.2.3 Actuator layer
This layer mainly includes controllable on-off and regulating valves as well as pumps and frequency converter sets. SED valves [10], CDR pumps [11] and Yamei vacuum self-priming pumps [12] are used. The specification of pumps are listed in Table 2.
Table 2: Pump specifications for control system
Pump type Quantity Purpose
Electromagnetic pump 2 Deliver LS from LS product tank to FOC system
Electromagnetic pump 2 Deliver LS from FOC system to CD or return it to purification system
Electromagnetic pump 2 Extract LS from CD and pump to the purification system for circulation
Vacuum self-priming pump 2 Extract pure water from CD during LS exchanging
All electromagnetic pumps operate in a “one operation and one standby” configuration. Real-time pump status monitoring is achieved through digital feedback signals (ready, fault, running) and analog feedback of actual operating frequency, which are compared against control signals. Pressure sensors downstream of pumps monitor hydraulic pressure to prevent overpressure, while upstream level interlocks safeguard against dry running. All valves come with ample spare parts and manual bypass valves. On-off valves provide “opened” or “closed” status feedback signals, while regulating valves offer analog feedback on opening degrees. Both pumps and valves have additional monitoring methods include local inverter displays and regular operator rounds, ensuring system stability.
2.2.4 Alarm and data management layer
The control system integrates a variety of alarm methods. In the control cabinet, an industrial sound and light alarm siren horn is installed. Once the system malfunctions, the alarm device will immediately trigger audiovisual alerts on-site operators to deal with it in a timely manner. The HMI displays anomalies through red flashing widgets and pop-up alerts for rapid identification. The physical emergency stop button ensures immediate system shutdown under extreme conditions.
In terms of data management, all data related to FOC are properly stored. The system employs a dual data storage architecture, with local hard disk storage enabling real-time analysis and troubleshooting, while synchronized Detector Control System (DCS) storage facilitates distributed access and multi-department data sharing. This integrated approach, coupled with web-based remote monitoring capabilities, ensures comprehensive data traceability.
3 Control logic
3.1 Architecture of the control logic
Building on the hardware architecture described in Section 2.2, drawing on the hierarchical control theory ISA-88 [13], the FOC control system employs a closed-loop control framework through three functional layers: detection, control, and monitoring & safety. This hierarchical design enables precise regulation of LS parameters while maintaining system stability during long-term operation.
  • Detection layer: signal acquisition and processing
    This layer interfaces with high-reliability sensors to collect real-time process parameters, employing a hybrid filtering strategy to mitigate noise. Arithmetic mean filtering reduces high-frequency fluctuation such as radar level gauge noise from 30 mm to 5 mm, while time hysteresis filter combines dual conditions of threshold and duration, avoiding false triggers caused by transient interference signals of Coriolis flow meters.
  • Control layer: algorithm implementation
    The control layer operates through four complementary control logic. Continuous control via Proportional-Integral-Derivative (PID) algorithms dynamically adjusts valve opening and pumping speed. State-machine-based sequential control enforces step-by-step execution, such as initiating LS exchanging only after tank level and temperature thresholds are met. The layer also integrates split-range control (which coordinates multiple actuators via a single controller) and selective control (which prioritizes critical parameters under competing constraints). Boolean logic interlocks provide additional protection, triggering actions such as shutting down pumps automatically if pressure goes beyond the safety limits. Detailed design and application scenarios will be introduced in Section 3.2.
  • Monitoring and safety layer: redundant safeguard
    Detailed specifications of the monitoring and safety layer have been elaborated in Section 2.2. This layer incorporates HMI monitoring and multiple alarm modes, enabling operators to monitor the process parameters and manually intervene in the control logic to trigger critical actions timely during system failures.
3.2 Typical control logic types and application scenarios
Multiple typical control logic functions guarantee the operation of the FOC system safely and efficiently, which includes: PID control, sequential control, safety interlock, and split-range & selective control logic.
  • PID control
    Following Ziegler-Nichols tuning rules [14], the PID control logic is primarily used for the closed-loop control of continuous variables such as flow rate. In the LS filling process, precise control of the liquid flow rate is critical. An excessive flow rate may undermine the structural safety of the CD, while an extremely slow flow will reduce the filling efficiency. The logical principle of PID control is described as follows:
[…]
    where $u(t)$ denotes the control output signal to actuators (e.g., regulating valves), $K_{p}$ is the proportional gain, which is directly scaled to the current deviation $e(t)$ (setpoint minus actual value) for immediate response, $T_{i}$ (integral time constant) accumulates past deviations to eliminate steady-state errors, ensuring long-term accuracy, while $T_{d}$ (derivative time constant) predicts system trends by reacting to the rate of change of the deviation, damping overshoots.
    Taking LS filling control as an example: the PID controller dynamically adjusts the flow through three coordinated actions, when actual flow falls below the setpoint, the proportional term immediately boosts output. The derivative term counters rapid flow increases by generating opposing signals based on the rate of deviation change, preventing overshoot. Simultaneously, the integral term continuously eliminates residual errors - during initial filling, strong proportional action dominates, while the integral term handles fine adjustments as flow nears the target, with derivative action maintaining stability throughout the transition.
  • Sequential control
    The sequential control logic employs a state-machine design to enforce strict stepwise execution of operations through condition-action chains, where each process phase has defined prerequisites and actions. For instance, LS exchanging initiation requires the storage tank to reach a preset level while meeting both the LS temperature specifications and calibration house pressure threshold, with the system only ramping up exchange speed after verifying stable low-flow filling to ensure detector safety.
  • Safety interlock
    The safety interlock logic mitigates accidents by sensing process parameters out of range and acting immediately. Risks during detector filling or operation, such as detector pressure exceeding the setpoint, abnormal LS exchanging flux or LS temperature, may damage the detector. The safety interlock logic following two basic principles: independence and fail-safety. For the independence principle, the safety logic operates independently of the normal control logic to avoid common-cause failures. For example, the detector calibration house pressure monitoring uses independent sensors in the safety interlock system, ensuring accurate monitoring and triggering of safety actions even if normal control sensors fail. The fail-safety principle means the system will start the safe action by default in case of failure. Critical equipment such as pumps and valves shut down automatically in case of CD pressure imbalance.
  • Split-range and selective control
    Split-range control plays an important role in flow control of the water filling process; a single controller output drives multiple actuators (valve group composed of two regulating valves) to work coherently and precisely. In selective control scenarios, the system implements priority-based logic that dynamically selects control commands from multiple controllers, where critical parameters (e.g., pressure and liquid level) are assigned higher precedence over secondary variables to ensure operational safety.
4 Test and performance
In alignment with different operational scenarios, multiple simulated condition experiments were conducted to replicate the real FOC working conditions.
4.1 Test with pure water system
Before pure water filling operations, several functional tests were performed in the FOC and water systems using the test pipeline (Fig. 1). The incoming water was divided into two streams [5]: one feeds directly to the WP, while the other undergoes ultrafiltration for CD use. The water flow is routed through control valves and test pipelines to the backwater, with vortex meters providing PID feedback for valve adjustment.
The test results (Fig. 4) demonstrate the rapid response and precise flow control of the FOC control system, with the WP ($39.3\ \text{m}^{3}/\text{h}$) and CD ($18.9\ \text{m}^{3}/\text{h}$) pipelines reaching the target flows in 2 minutes and maintaining stability within 0.5% uncertainty, meeting all design specifications.
(a) WP filling automatic control
(b) CD filling automatic control
Figure 4: WP and CD filling automatic control at test stage
4.2 Test with LS purification system
The joint test between the FOC system and the LS purification system successfully verified two operational modes: (1) LS transfer during LS filling (with OSIRIS radioactivity monitoring) and (2) continuous purification recirculation. The test confirmed reliable LS transfer, stable pump performance, and proper execution of sequential control logic and safety interlocks between systems, while maintaining all purity requirements throughout the process.
Figure 5: Liquid level changes during the test with LS purification system
During test, LS was transferred from the purification system to FOC tanks at varying flow rates, with automatic pump control triggered by tank level thresholds. Figure 5 shows the changes in liquid level at $4.5\ \text{m}^{3}/\text{h}$ during the transfer of LS between tanks. Although the comparison between differential pressure and radar level gauges confirms measurement accuracy (deviation < 0.2%), some limitations exist: (1) the differential pressure gauge exhibited a 0-0.3 m dead zones due to installation constraints; (2) The radar gauge demonstrates both a 0.2 m dead zone (for FOC horizontal tanks) and measurable response lag.
4.3 Internal test of FOC system
Overflowing test evaluated two key operational modes: (1) overflowing testing through LS transfer from storage to overflow tanks, verifying automatic replenishment when CD/overflow tank levels dropped below lower thresholds, and (2) overfill protection validation, where overflow valves automatically redirected excess LS back to storage after exceeding upper level limits. The tests also confirmed tank connectivity, selective control algorithms, and safety interlocks.
LS filling test utilized the overflow tank and temporary pipelines to validate automatic control performance. LS was pumped from the storage tank through a controlling valve, Coriolis flowmeter, and filter, with PID feedback maintaining the target flow rate ($1500$ L/h) with $0.5\%$ accuracy after a 3-minute stabilization period. This test not only verifies the accuracy matching of the flowmeter, but also visually demonstrates the dynamic response capability of the PID algorithm.
Internal test has proved the robust, safe, and accurate performance of the control system in overflowing and LS exchanging scenarios.
5 Conclusion
The FOC control system serves as the central platform for LS filling, overflowing, and circulating in the JUNO experiment. By integrating a high-stability PLC architecture, precision sensors (accuracy < 0.2% FS), and advanced control logic with PID algorithms, sequential control and safety interlocks, the system achieves closed-loop regulation of critical parameters with exceptional reliability. Test performances demonstrate the water/LS filling and the shock-free LS transfer with rapid stabilization in 3 minutes and flow rate control accuracy of 0.5%. Hardware redundancy such as modular PLCs, one-operational-one-standby pumps and software safeguards (e.g., dual HMIs, multi-type alarm mechanisms) ensure fault tolerance and anti-interference capability. These achievements establish a solid foundation for the safe and efficient filling of the 20-kiloton LS detector and its stable operation over 20-year lifespan.
Acknowledgments
This work is supported by NO.2023YFA1606100, National Key R&D Program of China.